In a circle whose radius is 15 inches, a chord is 12 inches from the center, what is the length of the chord: In a circle whose radius is 15 inches, a chord is 12 inches from the center, what is the length of the chord
Answer by Alan3354(1449)

This forms a right triangle with a hypotenuse of 15 (the radius) and one side of 12. The other side is 1/2 of the chord.
(C/2)^2 = 15^2 - 12^2
C/2 = sqrt(225 - 144)
C/2 = sqrt(81)
C/2 = 9
Chord = 18